OpenStudy (anonymous):

1/6x+4/5=7/10 helpp please pronto

OpenStudy (imtant):

Bring 4/5 to the right that will become 1/6x=(7/10)-4/5 1/6x=-1/10 x=-3/5..

OpenStudy (anonymous):

i got it thxs

OpenStudy (imtant):

Yw..